OpenStudy (anonymous):

Is 8/x+8=24 linear or non-linear and why?

OpenStudy (anonymous):

Is it \[\frac{ 8 }{ x+8 }=24\] or \[\frac{ 8 }{ x }+8=24\] but anyways best thing to do with be to graph it, and 1/x graph looks something like this...

OpenStudy (anonymous):

So what can we conclude from that information alone?

OpenStudy (anonymous):

That it is a linear equation?

OpenStudy (anonymous):

Mhm linear equations look like this